A 1200 kg car enters a round about (Area = 81 π) with velocity of 20 m/s, 1) find the Centripetal acceleration of the car. 2)find the required force to keep the car in it’s path


1
Expert's answer
2021-02-07T19:08:03-0500

"A=81\pi=2\pi r\to r=40.5\ (m)"


1)


"a=v^2/r=20^2/40.5=9.88\ (m/s^2)"


2)


"F=m\frac{v^2}{r}\to F=1200\cdot9.88=11852\ (N)" . Answer